Adult tickets for a play cost $4 and child tickets cost $1. If there were 23 people at a performance and the theater collected $80 from ticket sales, how many children attended the play?
Answer by stanbon(75887)

Adult tickets for a play cost $4 and child tickets cost $1. If there were 23 people at a performance and the theater collected $80 from ticket sales, how many children attended the play?
----------
Let number of children be "x" ; Value of their tickets is x dollars
Number of adult tickets is "23-x"; Value of these is 4(23-x)=(92-4x) dollars
EQUATION:
value + value = 80 dollars
x + 92-4x=80
-3x=-12
x=4 (number of children attending)
Cheers,
Stan H.